# Pareto Envelope Augmented with Reinforcement Learning: Multi-objective reinforcement learning-based approach for

### 19 Dec 2023 v2

@@ 5.3 Summary of the results obtained
In Section 5.1 we demonstrated the application of PEARL (and its constraints handling counterpart C-PEARL) on classical benchmarks. In Section 5.2 we apply these methods to the classical PWR fuel LP optimization problem in Two- and Three-objective settings. In these scenarios, we demonstrated the potential for PEARL to outperform legacy approaches in a small amount of computing samples, and emphasize the role and performance of the CL with distance- and rank-based techniques. Tables 12 and 13 summarize the performance of each algorithm studied on the benchmarks and against the NSGA approach, in the unconstrained and constrained cases, respectively. For each problem we assign ”1” if the performance was satisfying, while 0 if the performance was not. For clarity we also added a column ”Type for objectives” and ”Type for constraints”, which emphasize the method the agents are utilizing to traverse the search space. ”rank” means that the reward assigned to the solution generated is obtained by ranking the solution within a dynamically updated buffer, while ”distance” means that the pure magnitude of an objective is utilized instead. Overall, on all problems and both in the unconstrained and constrained cases, the PEARL-NdS surpasses all the PEARL methods as well as NSGA.
@@
PEARL-NdS (crowding) rank 5/5 2/2 Surpass
PEARL-e (cos) distance 5/5 ${}^{\star}$ Similar
PEARL-$\epsilon$ rank 5/5 1/2 Inferior
  • ${}^{\star}$this case could be improved by tweaking the $\alpha$ distribution.
@@
PEARL-NdS (niching) CL rank distance $-{}^{\star}$ 2/2 Surpass
PEARL-NdS (crowding2) rank rank - 2/2 Surpass
PEARL-NdS (niching2) ran rank 6/6 1/2 Similar
PEARL-e (cos) CL distance distance 6/6 - -
PEARL-$\epsilon$ CL rank distance 4/6 - -
  • ${}^{\star}$ Results marked with a ”-” were not shown in this paper.
@@ 6 Concluding remarks
In this study, we devised three variants of the Pareto Envelope Augmented with Reinforcement Learning (PEARL), namely PEARL-e, PEARL-$\epsilon$, and PEARL-NdS as detailed in Section 3.1. These variants were further expanded into C-PEARL in Section 3.2, incorporating a constraints satisfaction module. On top of that, thanks to the better performance of PEARL-NdS, we presented results for four sub-versions: PEARL-NdS (crowding) ,PEARL-NdS (crowding) in the unconstrained scenario, supplemented by PEARL-NdS (crowding2) and PEARL-NdS (niching2) in the constrained scenario. The former leverage a distance-based approach to solve the constraints, while the latter use a magnitude free approach by ranking solutions.
@@
In the context of the PWR problem, PEARL-NdS consistently outperforms classical SO-based approaches, particularly NSGA-III, across various multi-objective metrics, including the HV, as demonstrated in Tables 8, 10, 9, and 11. This advantage in HV is particularly noteworthy for core designers, as it enables more efficient exploration of the search space using PEARL-NdS. However, it’s important to note that future research should include a more comprehensive statistical analysis to confirm the significant superiority of the newly-developed PEARL approaches, in particular PEARL-NdS, over legacy methods like NSGA-III.
@@
Furthermore, we saw that on the PWR problems solved with PEARL-NdS, once the agents reach the trade-off area of the FOMs studied, they will preferentially optimize first the average enrichment, the $L_{C}$, and $F_{\Delta h}$ by order of simplicity in order to get higher rank and higher reward faster. As a result, even when two objective are nearly perfectly negatively correlated (which is the case of the average enrichment and $L_{C}$), the agents still had the potential to recover a high-quality Pareto front in case of the PEARL-NdS (crowding) and PEARL-NdS (niching) versions. However, as discussed in Section 5.2.2, we encountered a potential issue with the rank-based approaches PEARL-NdS (crowding2) and PEARL-NdS (niching2) in the constrained case. Once feasible solutions are discovered, the agents tend to prioritize the optimization of a single objective first before the others, resulting in slower learning and reduced sample efficiency. The application of the distance-based approaches PEARL-NdS (crowding) and PEARL-NdS (niching) to handle constraints introduces a distinct disparity in rewards between feasible and infeasible solutions. This disparity effectively mitigates the aforementioned effect and broadens the range of solutions found along the Pareto front.
To deal with constraints, we separated the problem between finding feasible solutions and finding the optimum Pareto envelope afterward (see Section 3.2). There, all constraints were solved at once. A natural enhancement of the method is to create multiple sub-problems combining different constraints together as in [19]. This study is the focus of our ongoing efforts.
@@
Lastly, replacing the average enrichment minimization by minimizing the BOC exposure might be a preferable target. Indeed, first, it compromises with $L_{C}$ [48] for the current cycle as it reduces the amount of energy extracted from the fuel in the cycle allowing more energy for the following one. The latter presents economic benefits when the end goal is multi-cycle optimization, which is the typical next step in PWR design [5, 48]. Future research could expand the application of PEARL to such setting.
@@ Acknowledgement
Appendix A Appendix: Classical benchmarks test suite
A.1 Unconstrained benchmarks
In this appendix, we present graphical results detailing the performance of various versions of PEARL on the classical unconstrained benchmarks discussed in Section 4.2.1. We previously examined these benchmarks using HV and $I_{C}$ metrics in Section 5.1.1. Our initial focus is on exploring the impact of ray tracing with PEARL-e, where we vary $\alpha$ for the Dirichlet distribution. The results are visually presented in Figure 6.
Figure 6: Results for PEARL-e with different $\alpha$ (from left to right) $\lambda$ $=0$, $\alpha$$=[0.5,0.5,0.5]$, $\alpha$$=[1.0,1.0,1.0]$, and $\alpha$$=[10,10,10]$ applied to the benchmarks (from top to bottom) dtlz2, dtlz4, dtlz5, dtlz6, dtlz7. In red, the pareto equivalent solutions generated by the algorithms. In black, pre-defined Pareto-optimal points.
First, let’s examine the impact of the absence of a uniformity factor (leftmost case), which consistently underperforms in comparison to the other cases, except for dtlz7, as we previously detailed in Table 2. Focusing specifically on dtlz2, where the Pareto surface is uniformly distributed across the objective space, we observe some interesting patterns. When rays are drawn at the center of the hyper-triangle (case $[10,10,10]$), the resulting solutions (represented by red dots) are densely clustered in the center. In contrast, for benchmarks where samples are drawn from preferential edges (case $[0.999,2.0,2.0]$), the red dots tend to cluster along those edges. Furthermore, for dtlz2, we notice that the case $[10,10,10]$ produces a higher number of optimal solutions but with a lower HV compared to the $[1,1,1]$ case. This could be attributed to the solutions found being overly localized in the former case. On top of that, for benchmarks where the Pareto front is sparse, the distribution matters more (e.g., dtlz4 and dtlz5). In cases where the Pareto front is sparse, the distribution of rays becomes crucial, as seen in dtlz4 and dtlz5. The $[10,10,10]$ case generates a higher number of solutions because it effectively captures the sparse nodes in the center of the Pareto front surface. In contrast, the $[0.999,2.0,2.0]$ case only captures the Pareto-optimal points along an edge.
While PEARL-e proves to be an efficient method, it does accentuate the challenge of potentially introducing computational overhead to adapt the rays to the specific problem at hand. This complexity may arise, particularly when the Pareto front is unknown. However, when a designer possesses some knowledge of the Pareto surface, fine-tuning the Dirichlet distribution with $\alpha$ could yield significant advantages.
Figure 7: Results of different versions of PEARL (from left to right) PEARL-NdS (crowding), PEARL-NdS (niching), PEARL-$\epsilon$, and classical legacy approaches NSGA-II/III applied to the benchmarks (from top to bottom) dtlz2, dtlz4, dtlz5, dtlz6, dtlz7. In red, the pareto equivalent solutions generated by the algorithms. In black, pre-defined Pareto-optimal points.
Conversely, we see on Figure 7 that the PEARL-NdS and PEARL-$\epsilon$ systematically found Pareto-optimal solutions without the need for any assumptions regarding ray distributions tailored to a specific front. This flexibility is highly beneficial for real-world practitioners who seek to utilize these tools without delving into problem-specific tuning. However, it’s worth noting that, in all cases for dtlz4, PEARL couldn’t capture points in the center of the front (although it excelled at identifying points on the edges) where legacy approaches succeeded. This specific challenge might warrant further targeted tuning for this problem.
A.2 Constrained benchmarks
In this appendix, we provide graphical results about the performance of the different versions of C-PEARL on the classical constrained benchmarks mentioned in Section 4.2.1 for which we studied the HV and $I_{C}$ metrics in Section 5.1.2. The results are displayed in Figure 8.
PEARL-based approaches consistently discover diverse and widely distributed Pareto-optimal points, indicated by the red dots in the figure. Notably, the assumption-free versions PEARL-NdS and niching2 outperform PEARL-e, which could be attributed to the arbitrary choice of $\alpha$$=[10,10,10]$. This highlights the challenge of fine-tuning $\alpha$ in PEARL-e.
For c3dtlz4, the absence of points in the center, except for niching2, mirrors the unconstrained case in dtlz4. This similarity arises because the front of this problem remains largely the same, with only the addition of constraints. The lack of visible red dots for NSGA-III on the cptx problems (rows 3 to 6) results from the limited number of Pareto-equivalent solutions found, with some cases having only two different solutions, as exemplified by ctp1.
The seemingly modest performance of ctp3 and ctp4 in Tables 5 and 6 is explained by the fact that the Pareto front reduces to a single point for both problems, as elaborated in Section 4.2.1. Nevertheless, the figure clearly illustrates that, for most PEARL algorithms, we obtain superior and more widely dispersed solutions compared to the legacy-based approach.
Figure 8: Result of different algorithms (from left to right) PEARL-e, PEARL-NdS (crowding) with CL, PEARL-NdS (niching2), PEARL-$\epsilon$, and NSGAIII applied to the benchmarks (from top to bottom) c2dtlz2, c3dtlz4, ctp1, ctp2, ctp3, ctp4. In red, the pareto equivalent solutions generated by the algorithms. In black, pre-defined Pareto-optimal points.

### 15 Mar 2024 v3

@@ 5.3 Summary of the results obtained
In Section 5.1 we demonstrated the application of PEARL (and its constraints handling counterpart C-PEARL) on classical benchmarks. In Section 5.2 we apply these methods to the classical PWR fuel LP optimization problem in Two- and Three-objective settings. In these scenarios, we demonstrated the potential for PEARL to outperform legacy approaches in a small amount of computing samples, and emphasize the role and performance of the CL with distance- and rank-based techniques. Tables 18 and 19 summarize the performance of each algorithm studied on the benchmarks and against the NSGA (best algorithm out of the SO-based methods) approach, in the unconstrained and constrained cases, respectively. For each problem we assign ”1” if the performance was satisfying, while 0 if the performance was not. For clarity we also added a column ”Type for objectives” and ”Type for constraints”, which emphasize the method the agents are utilizing to traverse the search space. ”rank” means that the reward assigned to the solution generated is obtained by ranking the solution within a dynamically updated buffer, while ”distance” means that the pure magnitude of an objective is utilized instead. Overall, on all problems and both in the unconstrained and constrained cases, the PEARL-NdS surpasses all the PEARL methods as well as NSGA.
@@
PEARL-NdS (niching) rank 7/7 $-{}^{\star}$ ${}^{\star\star}$
PEARL-NdS (crowding) rank 7/7 2/2 Surpass
PEARL-e (cos) distance 7/7 ${}^{\star\star\star}$ Inferior
PEARL-$\epsilon$ rank 7/7 1/2 Inferior
  • ${}^{\star}$ Results marked with a ”-” were not shown in this paper.
  • ${}^{\star\star}$this assumes that both (niching) and (crowding) have similar performance for the PWR cases.
  • ${}^{\star\star\star}$this case could be improved by tweaking the $\alpha$ distribution.
@@
PEARL-NdS (niching2) rank rank 8/8 2/2 Surpass
PEARL-NdS (crowding2) rank rank 8/8 2/2 Surpass
PEARL-e (cos) CL distance distance 8/8 - -
PEARL-$\epsilon$ CL rank distance 6/8 - -
@@ 6 Concluding remarks
In this study, we devised three variants of the Pareto Envelope Augmented with Reinforcement Learning (PEARL), namely PEARL-e, PEARL-$\epsilon$, and PEARL-NdS as detailed in Section 3.1. These variants were further expanded into C-PEARL in Section 3.2, incorporating a constraints satisfaction module. On top of that, thanks to the better performance of PEARL-NdS, we presented results for four sub-versions in Section 5.2.2: PEARL-NdS (crowding) ,PEARL-NdS (crowding) in the unconstrained scenario, supplemented by PEARL-NdS (crowding2) and PEARL-NdS (niching2) in the constrained scenario. The former leverage a distance-based approach to solve the constraints, while the latter use a magnitude-free approach by ranking solutions.
@@
In the context of the benchmarks, PEARL systematically outperforms TS and SA, but does not always outperform NSGA. However, even if in average the HV could be higher for NSGA in the unconstrained cases (e.g., for dtlz1 and dtlz3 in Section 5.1.1), there is never a statistically significant advantage. On the other hand, when PEARL is superior in HV, there is such advantage most of the time, especially in the constrained cases. This demonstrate the potential of PEARL to perform well for various types of problem and topologies of the Pareto front.
@@
In the context of the PWR problem, PEARL-NdS consistently outperforms classical SO-based approaches, particularly NSGA, SA, and TS, across various multi-objective metrics, including the HV, as demonstrated in Tables 14, 16, 15, and 17. This advantage in HV is particularly noteworthy for core designers, as it enables more efficient exploration of the search space using PEARL-NdS. However, it’s important to note that future research should include a more comprehensive statistical analysis to confirm the significant superiority of the newly-developed PEARL approaches, in particular PEARL-NdS, over legacy methods like NSGA, SA, and TS.
Additionally, we found similar conclusion than observed in our comparison between RL and SO-based approaches in single-objective settings in another research [53, 54]. There, we demonstrated that the PWR problems are more amenable to global search (i.e., large changes in the input candidates from one step to the next) in the beginning to find high-quality solutions rapidly, and then local search (i.e., small changes in the input candidates from one step to the next) to be able to observe continuous improvement over time. The policy in RL, which adapts over time and hence behaves in this way could explain its superiority in multi-objective optimization as well against NSGA-II and NSGA-III, which are only global-search algorithms, and SA and TS, which are only local-search algorithms.
@@
Furthermore, we saw that on the PWR problems solved with PEARL-NdS, once the agents reach the trade-off area of the FOMs studied, they will preferentially optimize first the average enrichment, the $L_{C}$, and $F_{\Delta h}$ by order of simplicity in order to get higher rank and higher reward faster. As a result, even when two objectivse are nearly perfectly negatively correlated (which is the case of the average enrichment and $L_{C}$), the agents still had the potential to recover a high-quality Pareto front in case of the PEARL-NdS (crowding) and PEARL-NdS (niching) versions. However, as discussed in Section 5.2.2, we encountered a potential issue with the rank-based approaches PEARL-NdS (crowding2) and PEARL-NdS (niching2) in the constrained case. Once feasible solutions are discovered, the agents tend to prioritize the optimization of a single objective first before the others, resulting in slower learning and reduced sample efficiency. The application of the distance-based approaches PEARL-NdS (crowding) and PEARL-NdS (niching) to handle constraints introduces a distinct disparity in rewards between feasible and infeasible solutions. This disparity effectively mitigates the aforementioned effect and broadens the range of solutions found along the Pareto front. To deal with constraints, we separated the problem between finding feasible solutions and finding the optimum Pareto envelope afterward (see Section 3.2). There, all constraints were solved at once. A natural enhancement of the method is to create multiple sub-problems combining different constraints together as in [19]. This study is the focus of our ongoing efforts.
@@
Lastly, replacing the average enrichment minimization by minimizing the BOC exposure might be a preferable target. Indeed, it compromises with $L_{C}$ [51] for the current cycle as it reduces the amount of energy extracted from the fuel in the cycle allowing more energy for the following one. The latter presents economic benefits when the end goal is multi-cycle optimization, which is the typical next step in PWR design [5, 51]. Future research could expand the application of PEARL to such setting.
@@
Data availability
PEARL will be a part of the open-source NEORL project (Neuro Evolution optimization with Reinforcement Learning) [60], which includes a set of implementations of hybrid algorithms combining neural networks and evolutionary computation. The documentation for the algorithms in NEORL and examples can be accessed via this link https://neorl.readthedocs.io/en/latest/modules/modules.html.